OpenStudy (notsmartatall):

solve 5x^3-6x^2-4x-8=0

OpenStudy (mathlegend):

I would factor by grouping... what can you take out of this... what do they have in common? \[5x ^{3}-6x ^{2}\]

OpenStudy (mathlegend):

@notsmartatall ?

OpenStudy (notsmartatall):

nun of them are in common

OpenStudy (mathlegend):

\[x ^{2}(5x-6)\]

OpenStudy (mathlegend):

Do you understand how I took out 2 x variables? X^3 + X^2 They both have 2 x variables... so I took 2 out @notsmartatall

OpenStudy (notsmartatall):

ohhh ya that makes sense now

OpenStudy (mathlegend):

So what about (-4x-8) What can you take out of this? What is the highest number that you can divide both of the numbers by?

OpenStudy (notsmartatall):

-4

OpenStudy (mathlegend):

-4(x+2) Right... so now we have another part

OpenStudy (mathlegend):

\[x ^{2}(5x-6)-4(x+2)\]

OpenStudy (mathlegend):

So you understand I just lined everything up right here @notsmartatall ?

OpenStudy (notsmartatall):

ya ok i get it

OpenStudy (mathlegend):

\[(x ^{2}-4)(5x-6)(x+2) = 0\]

OpenStudy (mathlegend):

So here I pulled down what ever was in parenthesis already. Then what was on the outside of the parenthesis I put them together in the first set.

OpenStudy (mathlegend):

\[x ^{2}-4 = 0 \]

OpenStudy (mathlegend):

We must solve all of them one by one now... so that one is first. Can you solve for x?

OpenStudy (notsmartatall):

x=2

OpenStudy (mathlegend):

Yes!

OpenStudy (mathlegend):

So that is one answer... now we move onto the next..

OpenStudy (mathlegend):

5x-6 = 0 Solve for x

OpenStudy (mathlegend):

x+2 = 0 Solve for x

OpenStudy (notsmartatall):

x=6/5

OpenStudy (mathlegend):

Good!

OpenStudy (notsmartatall):

x=-2

OpenStudy (mathlegend):

Good!

OpenStudy (mathlegend):

x = 2, -2, 6/5

OpenStudy (mathlegend):

So you are very smart!

OpenStudy (notsmartatall):

ahhaha thank you!!